OpenStudy (anonymous):

Use a proportion to solve. 1. 75% of what number is 66?

OpenStudy (anonymous):

Star let me see how u would set up your equation?

OpenStudy (katrinakaif):

Uh-oh

OpenStudy (anonymous):

Wanted star to set up the equation :D

OpenStudy (anonymous):

well i know that proportion is showing 2 equal ratios

OpenStudy (katrinakaif):

Use what I taught you.

hero (hero):

Katrina taught somebody something?

OpenStudy (anonymous):

75 of n=66 75x66=n

OpenStudy (anonymous):

I find Katrina very clear

OpenStudy (katrinakaif):

Come on star, break this down as I taught you too.

hero (hero):

is/of = %/100 66/x = 75/100

OpenStudy (anonymous):

hmm.. 66/75 = 0.88 0.88 * 100 = 88

OpenStudy (katrinakaif):

75% converts to?

OpenStudy (katrinakaif):

.75

OpenStudy (anonymous):

Guys wait you are all confusing star

hero (hero):

66/x = 3/4

OpenStudy (katrinakaif):

I guess Hero wants to explain. He shall assist you.

OpenStudy (anonymous):

Did I explain it properly?

hero (hero):

Don't look at anyone else's steps, just look at mine :P

OpenStudy (anonymous):

, 75 and 66 is .66 so

OpenStudy (katrinakaif):

66 remains 66 because it is not 66%. If it was a percentage, then you will use the decimal.

OpenStudy (anonymous):

ok

hero (hero):

It says to use a proportion. I provided the proper set up for this.

OpenStudy (katrinakaif):

Well then, provide the steps and help her.

hero (hero):

I did

OpenStudy (anonymous):

wait can u write all the steps down so i can write it down and study it

OpenStudy (anonymous):

so i can memorize

OpenStudy (katrinakaif):

Don't memorize Mathematics. Comprehend it.

OpenStudy (anonymous):

well teach me hero

OpenStudy (anonymous):

Katrina is correct =)

OpenStudy (katrinakaif):

as always =P

hero (hero):

The question asks: 75% of what number is 66? is/of = %/100 66/x = 75/100 66/x = 3/4 3*x = 66*4 3x = 264 x = 264/3 x = 88 Therefore, 75% of 88 is 66 Pippa, this truly is the proper solution given what is asked of the original question. Proportion is an equation with the form: a/b = c/d

OpenStudy (anonymous):

24 is 96% of what number?

OpenStudy (anonymous):

go

hero (hero):

I'm going to bed. I'm not doing this again, lol

OpenStudy (anonymous):

pippa can u explan it step by step?

OpenStudy (katrinakaif):

Star, try first.

OpenStudy (anonymous):

ok well 24n=96% and 96/n=24 so 96/24=4and244=96% so the answer is 6% of 96%

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

correction 24 is 6% of 96%

OpenStudy (katrinakaif):

Star...

OpenStudy (anonymous):

wat is it wrong agian?

OpenStudy (katrinakaif):

24 is 96% of what number? What is your expression here?

OpenStudy (anonymous):

24 is n of 96

OpenStudy (anonymous):

i thought i was right!!??

OpenStudy (katrinakaif):

96% converts to what?

OpenStudy (anonymous):

.96

OpenStudy (katrinakaif):

Good, so now your eequation is: 24 is n of .96

OpenStudy (katrinakaif):

"Is" converts to?

OpenStudy (anonymous):

um =

OpenStudy (anonymous):

%

OpenStudy (katrinakaif):

Great. Now our equation is: 24 = n of .96

OpenStudy (katrinakaif):

What does "of" convert to?

OpenStudy (anonymous):

%

OpenStudy (katrinakaif):

What does OF convert to Star?

OpenStudy (anonymous):

%

OpenStudy (katrinakaif):

"OF" does NOT convert to a percentage sign!!!

OpenStudy (katrinakaif):

Scroll up, view and refer to your past questions and my explanations on them!!!

OpenStudy (anonymous):

x

OpenStudy (katrinakaif):

Multiplication. Now our eqaution is??

OpenStudy (anonymous):

no i put x in cuz i meant mutliplycation i did not know how to put it

OpenStudy (katrinakaif):

Now, we found all our components. Our equation is?

OpenStudy (anonymous):

our equation is 24=n *.96

OpenStudy (anonymous):

i think thats times

OpenStudy (katrinakaif):

Yay! Good!

OpenStudy (katrinakaif):

Now lets solve. 24 = n x .96 The "n x .96" can be easily written as ".96n" Do you understand why?

OpenStudy (anonymous):

yes cuz when there togeter like that it is *(multiplycation)

OpenStudy (katrinakaif):

Good!! So our final equation is: 24 = .96n How can we get "n" alone?

OpenStudy (anonymous):

24 is 48% of 96%

OpenStudy (katrinakaif):

Answer my question..the answer you got is inaccurate.

OpenStudy (anonymous):

wat do u mean?

OpenStudy (katrinakaif):

24 = .96n Our goal is to get "n" alone. To do so, we have to divide .96 from both sides When we do so, we get: 24/.96 = .96n/.96 .96n / .96 gives us simply "n" 24/.96 give us?

OpenStudy (anonymous):

n=50%

OpenStudy (katrinakaif):

24 divided by .96 is???

OpenStudy (anonymous):

25

OpenStudy (katrinakaif):

Do you understand how to do this whole thing?

OpenStudy (anonymous):

yea just about thank u

OpenStudy (katrinakaif):

Anytime =)

OpenStudy (katrinakaif):

25 is correct by the by